40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Dec 12 17:22:07 2013, Steve, Update, General, rack power supplies checked 8x
    Reply  Fri Dec 13 16:34:20 2013, Steve, Update, VAC, PS will be replaced next week hotPSvacRack.jpgfan1X8.jpg
       Reply  Fri Dec 20 20:04:01 2013, Steve, Update, VAC, power supply replaced with a short vent 4hrPumpdown.pngpumpdownAfterHickup.pngPSpumpdown.png
          Reply  Fri Dec 20 23:00:41 2013, Koji, Update, VAC, power supply replaced with a short vent 
             Reply  Sat Dec 21 10:53:35 2013, Steve, Update, VAC, power supply replaced with a short vent-pumpdown completed 20hrsVacNormal.png
          Reply  Fri Jan 3 11:18:41 2014, Steve, Summary, VAC, power supply replaced with a short vent rebootVENT.png
    Reply  Thu Feb 22 10:49:11 2018, Steve, Update, Electronics, rack power supplies checked 8x
Message ID: 9508     Entry time: Fri Dec 20 23:00:41 2013     In reply to: 9506     Reply to this: 9510
Author: Koji 
Type: Update 
Category: VAC 
Subject: power supply replaced with a short vent 

I'm leaving the 40m now. IFO is aligned. Everything look good.

- The main volume P1=5e-4, CC1=1.4e-5 is still pumped by TP1 and TP2

- RGA P4<0e-4, CC4 2.1e-7, is pumped by TP3

- The annuluses are isolated.

- RP1/2/3 are off.

ELOG V3.1.3-